46 meters to kilometers
Answer by glabow(165)

The question is: How many meters are there in a kilometer? Kilo means 1000, so there are 1000 meters in a kilometer.
Now how much of a kilometer would 500 meters be? If you said 1/2 you would be right. That's because 500 meters/1000 meters = 1/2 = 0.5
How much of a kilometer would 1 meter be? Wouldn't it be 1/1000 meter? Yes.
So how much of a kilometer would 46 meters be. It looks like 46/1000 is the answer.
Of course, dividing by 1000 is the same as shifting the decimal point three places to the left.
46/1000 = 0.046
So, 0.046 kilometer = 46 meters.